OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 5x3 – 35x2 + 60x. Provide the factors.

OpenStudy (anonymous):

First factor out 5x then factor the remaining terms 5x(x^2-7x+12) 5x(x-3)(x-4)

OpenStudy (anonymous):

could you go in more detail and explanation

OpenStudy (anonymous):

I factored out 5x from the leaving (x^2-7x+12) which I factor again

OpenStudy (anonymous):

5x3-35x2+60x=x(5x2-35x+60) ax2+bx+c=a(x-x1)(x-x2) x1 and x2 are roots 5x2-35x+60, roots x1= 4 and x2=3 and a=5 5x2-35x+60= 5(x-4)(x-3) 5x3-35x+60x=x(5x2-35x+60)=x.5(x-4)(x-3)=5x(x-4)(x-3)